The reaction A+B rightarrow C+D rate=k[A][B]^2 ha
Answer #1

given that rate, r= [A] [B]2

A is halved and B is tripled New [A]=0.5[A] and new [B]= 3 [B}

new r= [0.5A] [3B]2= 9*0.5*[A] [B]2= 4.5 [A] [B]2 = 4.5*0.03= 0.135 M/s

b) A is tripled and B is halved

new r= k[3A] [0.5B]2 =3*0.25*[A] [B]2 =0.75 [A] [B]2 =0.75*0.03M/s =0.0225 M/s